OpenStudy (anonymous):

can you guys help me figure out the laplace transform of (t^2)sin(at) where a is a constant

OpenStudy (anonymous):

If I see this right: \[F(s)=L(f(t))=\int\limits_0^{\infty}e^{-st}f(t)dt\] Is this what you are trying to use?

OpenStudy (anonymous):

ya, thats it, i just can't figure out how to integrate that